The distribution of annual earnings of all bank tellers with five years of experience is skewed negatively. This distribution has a mean of Birr 15,000 and a standard deviation of Birr 2000. If we draw a random sample of 30 tellers, what is the probability that their earnings will average more than Birr 15,750 annually?
Answer by stanbon(75887)

The distribution of annual earnings of all bank tellers with five years of experience is skewed negatively.
This distribution has a mean of Birr 15,000 and a standard deviation of Birr 2000. If we draw a random sample of 30 tellers, what is the probability that their earnings will average more than Birr 15,750 annually?
z(15750) = (15750-15000)/(2000/sqrt(30)) = 750*sqrt(30)/2000 = 375*sqrt(30)/1000 = 0.375*sqrt(30) = 2.0540
---
P(x-bar > 15,750) = P(z > 2.0540) = normalcdf(2.0540,100) = 0.02
-----
Cheers,
Stan H.
----------